My problem is solve the equation for the variable A. As well as show how to get to that solution.
(A+B)/C=D

Answer by DWL(56)

You need to factor the solution for A
(A+B)/C=D
Multiply both sides by C: A+B = DC
Then subtract B from both sides to get: A = DC-B
Answer: A = DC-B